40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 1327     Entry time: Wed Mar 2 01:57:07 2011
Author: Zach 
Type: Laser 
Category: GYRO 
Subject: gyro locked with new RFPD 

 [Alastair, Zach]

We finally got all the parts together to properly mount the RFPD board to the box, so we did that today. We aligned the cavity mirrors and injected the primary beam in the CCW direction. I used the mode matching solution that I came up with last week, but there is a problem with it in that one of the mirrors has to occupy the same space as a turning mirror for the other beam. I will have to come up with a different solution, but I put the lenses as close to the right place as possible for right now.

Another issue is that---like before---it is going to be very tough to have the faraday isolators so close to the cavities while still having the beam small enough to fit through the center. We have a little more space on the table now, so we could conceivably move them further upstream in the MMT, but this makes the REFL extraction a bit messier. I think someone said that getting new enclosures with larger apertures for the polarizers flanking the rotator is just as difficult as getting a whole new FI, but we may have to figure a way to get the beam through, or else deal with clipping. Since there is no CW beam yet and I have to fix the mode matching of the CCW beam, I have not installed either faraday yet.

In any case, once I got some good TEM00 flashes, I distributed the LO signal from the Tektronix FG as before:

  • Coupler: 23 dBm IN from FG, ~23 dBm OUT to EOM through resonant circuit (which has a resonant gain of ~20 dB), 3 dBm CPL out to PDH mixer LOs
  • Splitter: 3 dBm IN from CPL out of coupler, 0 dBm out to each PDH mixer. Only one is in use now, so I terminated output 2.

I then installed RFPD S/N 01 and put the ~80 mW reflection from the cavity onto it. The DC output was consistent with the design, and dips were evident in conjunction with the spikes in the TRANS signal. I then connected the RF out to the RF IN of the mixer and swept the cavity to see the error signal. At first, I thought what i was getting was some distorted junk, but then I realized that the signal was just so huge that it was saturating the INPUT MON. Turning down the optical power resulted in the classic PDH signal sweep. Nevertheless, I was able to lock the cavity with the full 80 mW, though some UGF instability was evident at this level, so I had to turn it down to about 40 mW.

With the cavity locked, I iteratively adjusted both the cavity mirrors and the input steering mirrors to maximize transmission. I was able to get about 50% after quite a bit of trying, but I'm not convinced I can't do better even with the suboptimal mode matching.

Here are a couple shots of the PD doing its thing. The backplate is not on the box because I did not have a d-sub breakout board to power the thing.

newpd1.pngnewpd2.png

The plan for the immediate future is the following:

  1. Wait on and install remaining parts of RFPD S/N 02
    1. LM317 for bias -- none in Downs, ordered, on way(?)
    2. Cougar AP1053 -- deliberation in progress. We may be able to bypass the Cougar for the second PD temporarily using trickery.
  2. Finish tweaking cavity eigenmode, then pump cavity down
  3. Inject CW beam and install CW/CCW isolation optics
  4. Lock CW using PDH box S/N 2215 as-is.
  5. Measure gyro noise, compare to what is expected
  6. In parallel with above, build breadboard version of PDH2 servo and use it to lock the CCW beam. This will give us the real low-frequency improvement we want to report at the meeting
  7. Signal optimization, noise/loop characterization
ELOG V3.1.3-